OpenStudy (anonymous):

The product of 5 and b cubed

OpenStudy (diyadiya):

\[5 \times b^3=5b^3\]

OpenStudy (diyadiya):

\[(5b)^3=125b^3\]

OpenStudy (anonymous):

Thaannksss!!!